# Physics-Informed Teleconnection-Aware Transformer for Global Subseasonal-to-Seasonal Forecasting

### v2

@@ Abstract
Subseasonal-to-seasonal (S2S) forecasting, which predicts climate conditions from several weeks to months in advance, presents significant challenges due to the chaotic dynamics of atmospheric systems and complex interactions across multiple scales. Current approaches often fail to explicitly model underlying physical processes and teleconnections that are crucial at S2S timescales. We introduce TelePiT, a novel deep learning architecture that enhances global S2S forecasting through integrated multi-scale physics and teleconnection awareness. Our approach consists of three key components: (1) Spherical Harmonic Embedding, which accurately encodes global atmospheric variables onto spherical geometry; (2) Multi-Scale Physics-Informed Neural ODE, which explicitly captures atmospheric physical processes across multiple learnable frequency bands; (3) Teleconnection-Aware Transformer, which models critical global climate interactions through tactfully injecting teleconnection patterns into the self-attention. Extensive experiments demonstrate that TelePiT significantly outperforms state-of-the-art data-driven baselines and operational numerical weather prediction systems, with remarkable improvements for atmospheric variables including a 57.7% reduction in RMSE for 2-meter temperature compared to previous best models.
1 Introdution
@@
Subseasonal-to-seasonal (S2S) forecasting, predicting climate conditions from several weeks to months, occupies a critical frontier in atmospheric science robertson2015improving ; white2017potential ; zhang2025unleashing . By bridging the gap between short-range weather forecasts and seasonal climate projections, S2S outlooks are indispensable for proactive climate resilience and the efficient management of resources in agriculture, energy, and disaster preparedness white2017potential ; lyu2024autostf ; zhang2024irregular ; lyu2024nrformer . Compared to medium-range forecasting (up to 2 weeks) bi2023accurate ; lam2023learning , S2S forecasting encounters significantly greater challenges due to the inherent chaotic dynamics of atmospheric systems and intricate interactions across various scales of Earth’s climate system. Beyond initial atmospheric conditions, accurate S2S forecasting requires consideration of additional complex factors, such as more complicated and highly coupled physical processes, teleconnections, and boundary interactions. These complexities often render existing medium-range forecasting models inadequate for S2S forecasting, leading to a significant decline in predictive skill at extended lead times nathaniel2024chaosbench .
@@
Recent models have delivered promising improvements in both regional and global S2S forecasting tasks. While regional models mouatadid2023subseasonalclimateusa ; he2022learning ; hwang2019improving have significantly advanced localized forecasting capabilities, they often overlook the crucial influence of long-range correlations in atmospheric conditions, which become increasingly significant at S2S timescales. To achieve global forecasting, Chen et al.chen2024machine present FuXi-S2S, an encoder-decoder framework trained on global ERA5 reanalysis data hersbach2020era5 , notably enhancing global precipitation forecast. Furthermore, Liu et al. liucirt introduce CirT, a geometry-inspired transformer that effectively models Earth’s spherical structure for S2S forecasting.
@@
However, current methods still exhibit critical gaps. Firstly, existing works overlook explicit modeling of the underlying physical processes that dictate multi-scale climate dynamics. This oversight stems from reliance on fixed-frequency modeling and insufficient integration of physics-informed mechanisms, which are essential for accurate modeling of fundamental atmospheric processes such as advection, diffusion, and external forcing raissi2019physics ; karniadakis2021physics ; beucler2021enforcing . Secondly, many models fail to explicitly represent essential teleconnections (i.e.,  significant long-range correlations between distant regions), exemplified by climate modes such as North Atlantic Oscillation and Madden–Julian Oscillation. These teleconnections substantially influence global atmospheric conditions and play pivotal roles in determining regional and global climate variability on S2S scales trenberth1998atmospheric ; ham2019deep .
@@
2 Preliminaries
In this work, we represent the global weather condition on the day $\color[rgb]{0.171875,0.2421875,0.3125}t$ as $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{X}_{t}\in\mathbb{R}^{C\times H\times W}$, where $\color[rgb]{0.171875,0.2421875,0.3125}H$ and $\color[rgb]{0.171875,0.2421875,0.3125}W$ correspond to the spatial dimensions along latitude and longitude respectively, and $\color[rgb]{0.171875,0.2421875,0.3125}C$ denotes the number of weather variables (e.g., temperature, geopotential height). The latitude-longitude grid is defined as $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{G}\in\mathbb{R}^{H\times W}$, where each grid point $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{G}_{h,w}=(\theta_{h},\phi_{w})\in[-90^{\circ},90^{\circ}]\times[-180^{\circ},180^{\circ}]$.
S2S Forecasting
Following previous works chen2024machine ; liucirt , given the initial condition $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{X}_{t_{1}}$, the goal is to predict the mean fields of these variables over two future multi-week windows: weeks 3–4 (days 15–28 ) and weeks 5–6 (days 29–42) of the forecast period, as shown in the following:
$\mathcal{F}_{\Theta}(\mathbf{X}_{t_{1}})\rightarrow(\hat{\mathbf{Y}}_{t_{15}:t_{28}},\hat{\mathbf{Y}}_{t_{29}:t_{42}})$ (1)
where $\color[rgb]{0.171875,0.2421875,0.3125}\hat{\mathbf{Y}}_{t_{15}:t_{28}}\in\mathbb{R}^{C\times H\times W}$, and $\color[rgb]{0.171875,0.2421875,0.3125}\hat{\mathbf{Y}}_{t_{29}:t_{42}}\in\mathbb{R}^{C\times H\times W}$ denote the average value over weeks 3-4 and weeks 5-6 respectively. $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{F}_{\Theta}$ is forecasting model. S2S forecasting transcends the deterministic chaos limit of weather forecasting (about 2 weeks), thereby requiring advanced neural architecture to bridge the fundamental gap between chaotic weather dynamics and more predictable climate signals.
3 Methods
@@
3.1 Spherical Harmonic Embedding
To accurately represent global atmospheric variables on the sphere, we introduce a learnable spherical harmonic-inspired positional embedding. For a grid point at latitude $\color[rgb]{0.171875,0.2421875,0.3125}\theta_{i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and longitude $\color[rgb]{0.171875,0.2421875,0.3125}\phi_{j}\in[-\pi,\pi]$, we construct positional encodings using frequency-scaled sinusoidal functions:
$\mathbf{E}_{lat}(i,2k)=\sin\big{(}(k+1)\theta_{i}\big{)},\quad\mathbf{E}_{lat}(i,2k+1)=\cos\big{(}(k+1)\theta_{i}\big{)},$ (2)
$\mathbf{E}_{lon}(j,2q)=\sin\big{(}(q+1)\phi_{j}\big{)},\quad\mathbf{E}_{lon}(j,2q+1)=\cos\big{(}(q+1)\phi_{j}\big{)},$ (3)
where $\color[rgb]{0.171875,0.2421875,0.3125}k\in\{0,1,...,\frac{d_{lat}}{2}-1\}$ and $\color[rgb]{0.171875,0.2421875,0.3125}q\in\{0,1,...,\frac{d_{lon}}{2}-1\}$. These encodings are initialized as learnable parameters $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{E}_{lat}\in\mathbb{R}^{H\times d_{lat}}$ and $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{E}_{lon}\in\mathbb{R}^{W\times d_{lon}}$ where $\color[rgb]{0.171875,0.2421875,0.3125}d_{lat}+d_{lon}=D_{emb}$. The sinusoidal functions provide a natural way to represent the periodic structure of spherical coordinates, particularly for longitude which wraps around the globe atkinson2012spherical ; cohen2018spherical .
Figure 1: The framework of TelePiT for global subseasonal-to-seasonal forecasting.
We process the latitude dimension as a sequence of tokens while aggregating information along the longitude dimension. For each latitude $\color[rgb]{0.171875,0.2421875,0.3125}i$, we compute a zonal average: $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{u}_{i}=\frac{1}{W}\sum_{j=1}^{W}X_{i,j}\in\mathbb{R}^{C}$. This zonal averaging operation is physically justified as atmospheric patterns often exhibit stronger correlations along latitude bands, particularly for large-scale circulation features baldwin2003stratospheric ; kidston2015stratospheric . The features are then projected to a higher-dimensional latent space through a learnable transformation: $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{h}_{i}=\mathbf{W}\mathbf{u}_{i}+\mathbf{b}\in\mathbb{R}^{D_{emb}}$, where $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{W}\in\mathbb{R}^{D_{emb}\times C}$ and $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{b}\in\mathbb{R}^{D_{emb}}$ are learnable parameters.
@@ 3.1 Spherical Harmonic Embedding
$\mathbf{p}_{i}^{lat}=\mathbf{E}_{lat}(i,:)\in\mathbb{R}^{d_{lat}},\enspace\mathbf{p}^{lon}=\frac{1}{W}\sum_{j=1}^{W}\mathbf{E}_{lon}(j,:)\in\mathbb{R}^{d_{lon}},\enspace\mathbf{p}_{i}=[\mathbf{p}_{i}^{lat},\mathbf{p}^{lon}]\in\mathbb{R}^{D_{emb}}.$ (4)
@@
$\mathbf{z}_{i}=\mathbf{W}_{proj}(\mathbf{h}_{i}+\mathbf{p}_{i})+\mathbf{b}_{proj}\in\mathbb{R}^{D_{emb}},\enspace\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{H}]\in\mathbb{R}^{H\times D_{emb}}.$ (5)
@@
The theoretical foundation of this connects to spherical function representation theory. Mathematically, our zonal averaging and embedding method implements learnable approximation to spherical harmonic decomposition. Traditionally, a function $\color[rgb]{0.171875,0.2421875,0.3125}f(\theta,\phi)$ on sphere can be decomposed as: $\color[rgb]{0.171875,0.2421875,0.3125}f(\theta,\phi)=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}a_{\ell m}Y_{\ell}^{m}(\theta,\phi)$, where $\color[rgb]{0.171875,0.2421875,0.3125}Y_{\ell}^{m}(\theta,\phi)$ are the spherical harmonic functions and $\color[rgb]{0.171875,0.2421875,0.3125}a_{\ell m}$ are expansion coefficients atkinson2012spherical . In particular, when $\color[rgb]{0.171875,0.2421875,0.3125}m=0$, spherical harmonics reduce to zonal harmonics that depend solely on latitude $\color[rgb]{0.171875,0.2421875,0.3125}\theta$, which precisely corresponds to our zonal averaging operation. This approach extracts a sequence of length $\color[rgb]{0.171875,0.2421875,0.3125}H$ as input tokens for subsequent modules, preserving the inductive bias of spherical geometry while enabling the model to adaptively learn optimal representations for atmospheric variables.
3.2 Multi-Scale Physics-Informed Neural ODE
Learnable Multi-Scale Decomposition
@@ 3.2 Multi-Scale Physics-Informed Neural ODE
Atmospheric processes span multiple scales that interact in complex ways, creating a fundamental challenge for S2S forecasting. For example, when predicting Northern Hemisphere winter circulation, stratospheric polar vortex dynamics (planetary-scale phenomena) exhibit significant downward influence on tropospheric weather patterns (regional-scale outcomes) through scale-bridging mechanisms operating over 2-3 week baldwin2001stratospheric ; kidston2015stratospheric . These cross-scale interactions, also documented in observational studies butler2019predictability ; vitart2014evolution , embody critical physical processes that must be explicitly resolved to overcome current predictability limitations.
@@
To explicitly capture this inherent multi-scale structure, we implement an $\color[rgb]{0.171875,0.2421875,0.3125}L$-level learnable wavelet decomposition $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{W}$ operating along the latitude dimension of $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{Z}$. At each level, the current signal is split into a coarse approximation and a detail component, mirroring the discrete wavelet transform but with learnable filters. Formally, let $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{A}_{0}=\mathbf{Z}\in\mathbb{R}^{H\times D_{emb}}$ be the input at the finest scale. For each level $\color[rgb]{0.171875,0.2421875,0.3125}\ell=1,2,\cdots,L$, we apply a neural transform: $\color[rgb]{0.171875,0.2421875,0.3125}[\mathbf{A}_{\ell},\mathbf{D}_{\ell}]=\textit{split}\Big{(}\text{MLP}_{\ell}(\mathbf{A}_{\ell-1}),\text{dim}=2\Big{)}$, where $\color[rgb]{0.171875,0.2421875,0.3125}\text{MLP}_{\ell}:\mathbb{R}^{D_{emb}}\rightarrow\mathbb{R}^{2D_{emb}}$ is a two-layer perceptron with GELU activation, and the $\color[rgb]{0.171875,0.2421875,0.3125}\textit{split}(\cdot,\text{dim}=2)$ operation divides the output features into two equal halves. This yields an approximation $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{A}_{\ell}\in\mathbb{R}^{H\times D_{emb}}$ and a detail $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{D}_{\ell}\in\mathbb{R}^{H\times D_{emb}}$. The $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{A}_{\ell}$ carries the smoother (low-frequency) information, while $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{D}_{\ell}$ captures the variations (high-frequency details) removed at level $\color[rgb]{0.171875,0.2421875,0.3125}\ell$. After $\color[rgb]{0.171875,0.2421875,0.3125}L$ levels, we obtain a final coarse approximation $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{A}_{L}$ and detail components $\color[rgb]{0.171875,0.2421875,0.3125}\{\mathbf{D}_{\ell}\}_{\ell=1}^{L}$. For convenience, we denote the set of multi-scale features as: $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{X}=\{\mathbf{X}^{(0)},\mathbf{X}^{(1)},\dots,\mathbf{X}^{(L)}\}$, where $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{X}^{(0)}:=\mathbf{A}_{L}$ represents the lowest-frequency band, and $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{X}^{(\ell)}:=\mathbf{D}_{L+1-\ell}$ for $\color[rgb]{0.171875,0.2421875,0.3125}\ell=1,\cdots,L$ (ordering from low to high frequency). Each $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{X}^{(\ell)}\in\mathbb{R}^{H\times D_{emb}}$ maintains the same dimensions as the original embedding.
@@
This multi-scale decomposition aligns with the established understanding in atmospheric science that different physical processes dominate at different scales hoskins1985use , enabling our learnable approach to adaptively determine the optimal decomposition basis for S2S forecasting.
Physics-Informed Neural ODE

### v3

@@ Abstract.
Subseasonal-to-seasonal (S2S) forecasting, which predicts climate conditions from several weeks to months in advance, represents a critical frontier for agricultural planning, energy management, and disaster preparedness. However, it remains one of the most challenging problems in atmospheric science, due to the chaotic dynamics of atmospheric systems and complex interactions across multiple scales. Current approaches often fail to explicitly model underlying physical processes and teleconnections that are crucial at S2S timescales. We introduce TelePiT, a novel deep learning architecture that enhances global S2S forecasting through integrated multi-scale physics and teleconnection awareness. Our approach consists of three key components: (1) Spherical Harmonic Embedding, which accurately encodes global atmospheric variables onto spherical geometry; (2) Multi-Scale Physics-Informed Neural ODE, which explicitly captures atmospheric physical processes across multiple learnable frequency bands; (3) Teleconnection-Aware Transformer, which models critical global climate interactions through explicitly modeling teleconnection patterns into the self-attention. Extensive experiments demonstrate that TelePiT significantly outperforms state-of-the-art data-driven baselines and operational numerical weather prediction systems across all forecast horizons, marking a significant advance toward reliable S2S forecasting.
Global Subseasonal-to-Seasonal Forecasting, Climate Prediction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Subseasonal-to-seasonal (S2S) forecasting, predicting climate conditions from several weeks to months, occupies a critical frontier in atmospheric science (Robertson et al., 2015; White et al., 2017; Zhang et al., 2025). By bridging the gap between short-range weather forecasts and seasonal climate projections, S2S outlooks are indispensable for proactive climate resilience and the efficient management of resources in agriculture, energy, and disaster preparedness (White et al., 2017; Lyu et al., 2025b; Zhang et al., 2024; Lyu et al., 2025a). Compared to medium-range forecasting (up to 2 weeks) (Bi et al., 2023; Lam et al., 2023), S2S forecasting encounters significantly greater challenges due to the inherent chaotic dynamics of atmospheric systems and intricate interactions across various scales of Earth’s climate system. Beyond initial atmospheric conditions, accurate S2S forecasting requires consideration of additional complex factors, such as more complicated and highly coupled physical processes, teleconnections, and boundary interactions. These complexities often render existing medium-range forecasting models inadequate for S2S forecasting, leading to a significant decline in predictive skill at extended lead times (Nathaniel et al., 2024).
@@
Recent models have delivered promising improvements in both regional and global S2S forecasting tasks. While regional models (Mouatadid et al., 2023; He et al., 2022; Hwang et al., 2019) have significantly advanced localized forecasting capabilities, they often overlook the crucial influence of long-range correlations in atmospheric conditions, which become increasingly significant at S2S timescales. To achieve global forecasting, Chen et al.(Chen et al., 2024) present FuXi-S2S, an encoder-decoder framework trained on global ERA5 reanalysis data (Hersbach et al., 2020), notably enhancing global precipitation forecast. Furthermore, Liu et al. (Liu et al., 2025) introduce CirT, a geometry-inspired transformer that effectively models Earth’s spherical structure for S2S forecasting.
@@
However, current methods still exhibit critical research gaps. Firstly, existing works overlook explicit modeling of the underlying physical processes that dictate multi-scale climate dynamics. This oversight stems from reliance on fixed-frequency modeling and insufficient integration of physics-informed mechanisms, which are essential for accurate modeling of fundamental atmospheric processes such as advection, diffusion, and external forcing (Raissi et al., 2019; Karniadakis et al., 2021; Beucler et al., 2021). Secondly, many models fail to explicitly represent essential teleconnections (i.e.,  significant long-range correlations between distant regions), exemplified by climate modes such as North Atlantic Oscillation and Madden–Julian Oscillation. These teleconnections substantially influence global atmospheric conditions and play pivotal roles in determining regional and global climate variability on S2S scales (Trenberth, 1998; Ham et al., 2019).
@@
2. Preliminaries
In this work, we represent the global weather condition on the day $t$ as $\mathbf{X}_{t}\in\mathbb{R}^{C\times H\times W}$, where $H$ and $W$ correspond to the spatial dimensions along latitude and longitude respectively, and $C$ denotes the number of weather variables (e.g., temperature, geopotential height). The latitude-longitude grid is defined as $\mathcal{G}\in\mathbb{R}^{H\times W}$, where each grid point $\mathcal{G}_{h,w}=(\theta_{h},\phi_{w})\in[-90^{\circ},90^{\circ}]\times[-180^{\circ},180^{\circ}]$.
S2S Forecasting. Following previous works (Chen et al., 2024; Liu et al., 2025), given the initial condition $\mathbf{X}_{t_{1}}$, the goal is to predict the mean fields of these variables over two future multi-week windows: weeks 3–4 (days 15–28 ) and weeks 5–6 (days 29–42) of the forecast period, as shown in the following:
(1) $\mathcal{F}_{\Theta}(\mathbf{X}_{t_{1}})\rightarrow(\hat{\mathbf{Y}}_{t_{15}:t_{28}},\hat{\mathbf{Y}}_{t_{29}:t_{42}})$
where $\hat{\mathbf{Y}}_{t_{15}:t_{28}}\in\mathbb{R}^{C\times H\times W}$, and $\hat{\mathbf{Y}}_{t_{29}:t_{42}}\in\mathbb{R}^{C\times H\times W}$ denote the average value over weeks 3-4 and weeks 5-6 respectively. $\mathcal{F}_{\Theta}$ is forecasting model. S2S forecasting transcends the deterministic chaos limit of weather forecasting (2 weeks), thereby requiring advanced neural architecture to bridge the fundamental gap between chaotic weather dynamics and more predictable climate signals.
Figure 1. The framework of TelePiT for global subseasonal-to-seasonal forecasting.
3. Methods
@@
3.1. Spherical Harmonic Embedding
To accurately represent global atmospheric variables on the sphere, we introduce a learnable spherical harmonic-inspired positional embedding. For a grid point at latitude $\theta_{i}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ and longitude $\phi_{j}\in[-\pi,\pi]$, we construct positional encodings using frequency-scaled sinusoidal functions:
(2) $\mathbf{E}_{lat}(i,2k)=\sin\big{(}(k+1)\theta_{i}\big{)},\quad\mathbf{E}_{lat}(i,2k+1)=\cos\big{(}(k+1)\theta_{i}\big{)},$
(3) $\mathbf{E}_{lon}(j,2q)=\sin\big{(}(q+1)\phi_{j}\big{)},\quad\mathbf{E}_{lon}(j,2q+1)=\cos\big{(}(q+1)\phi_{j}\big{)},$
where $k\in\{0,1,...,\frac{d_{lat}}{2}-1\}$ and $q\in\{0,1,...,\frac{d_{lon}}{2}-1\}$. These encodings are initialized as learnable parameters $\mathbf{E}_{lat}\in\mathbb{R}^{H\times d_{lat}}$ and $\mathbf{E}_{lon}\in\mathbb{R}^{W\times d_{lon}}$ where $d_{lat}+d_{lon}=D_{emb}$. The sinusoidal functions provide a natural way to represent the periodic structure of spherical coordinates, particularly for longitude which wraps around the globe (Atkinson and Han, 2012; Cohen et al., 2018).
We process the latitude dimension as a sequence of tokens while aggregating information along the longitude dimension. For each latitude $i$, we compute a zonal average: $\mathbf{u}_{i}=\frac{1}{W}\sum_{j=1}^{W}X_{i,j}\in\mathbb{R}^{C}$. This zonal averaging operation is physically justified as atmospheric patterns often exhibit stronger correlations along latitude bands, particularly for large-scale circulation features (Baldwin et al., 2003; Kidston et al., 2015). The features are then projected to a higher-dimensional latent space through a learnable transformation: $\mathbf{h}_{i}=\mathbf{W}\mathbf{u}_{i}+\mathbf{b}\in\mathbb{R}^{D_{emb}}$, where $\mathbf{W}\in\mathbb{R}^{D_{emb}\times C}$ and $\mathbf{b}\in\mathbb{R}^{D_{emb}}$ are learnable parameters.
@@ 3.1. Spherical Harmonic Embedding
(4) $\mathbf{p}_{i}^{lat}=\mathbf{E}_{lat}(i,:)\in\mathbb{R}^{d_{lat}},\enspace\mathbf{p}^{lon}=\frac{1}{W}\sum_{j=1}^{W}\mathbf{E}_{lon}(j,:)\in\mathbb{R}^{d_{lon}},$
(5) $\mathbf{p}_{i}=[\mathbf{p}_{i}^{lat},\mathbf{p}^{lon}]\in\mathbb{R}^{D_{emb}}.$
@@
(6) $\mathbf{z}_{i}=\mathbf{W}_{proj}(\mathbf{h}_{i}+\mathbf{p}_{i})+\mathbf{b}_{proj}\in\mathbb{R}^{D_{emb}},$
(7) $\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{H}]\in\mathbb{R}^{H\times D_{emb}}.$
The theoretical foundation of this connects to spherical function representation theory. Mathematically, our zonal averaging and embedding method implements learnable approximation to spherical harmonic decomposition. Traditionally, a function $f(\theta,\phi)$ on sphere can be decomposed as:
(8) $f(\theta,\phi)=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}a_{\ell m}Y_{\ell}^{m}(\theta,\phi),$
@@
where $Y_{\ell}^{m}(\theta,\phi)$ are the spherical harmonic functions and $a_{\ell m}$ are expansion coefficients (Atkinson and Han, 2012). In particular, when $m=0$, spherical harmonics reduce to zonal harmonics that depend solely on latitude $\theta$, which precisely corresponds to our zonal averaging operation. This approach extracts a sequence of length $H$ as input tokens for subsequent modules, preserving the inductive bias of spherical geometry while enabling the model to adaptively learn optimal representations for atmospheric variables.
3.2. Multi-Scale Physics-Informed Neural ODE
3.2.1. Learnable Multi-Scale Decomposition
@@
Atmospheric processes span multiple scales that interact in complex ways, creating a fundamental challenge for S2S forecasting. For example, when predicting Northern Hemisphere winter circulation, stratospheric polar vortex dynamics (planetary-scale phenomena) exhibit significant downward influence on tropospheric weather patterns (regional-scale outcomes) through scale-bridging mechanisms operating over 2-3 week (Baldwin and Dunkerton, 2001; Kidston et al., 2015). These cross-scale interactions, also documented in observational studies (Butler et al., 2019; Vitart, 2014), embody critical physical processes that must be explicitly resolved to overcome current limitations.
@@
To explicitly capture this inherent multi-scale structure, we implement an $L$-level learnable wavelet decomposition $\mathcal{W}$ operating along the latitude dimension of $\mathbf{Z}$. At each level, the current signal is split into a coarse approximation and a detail component, mirroring the discrete wavelet transform but with learnable filters. Formally, let $\mathbf{A}_{0}=\mathbf{Z}\in\mathbb{R}^{H\times D_{emb}}$ be the input at the finest scale. For each level $\ell=1,2,\cdots,L$, we apply a neural transform:
(9) $\mathbf{A}_{\ell},\mathbf{D}_{\ell}]=\textit{split}\Big{(}\text{MLP}_{\ell}(\mathbf{A}_{\ell-1}),\text{dim}=2\Big{)},$
where $\text{MLP}_{\ell}:\mathbb{R}^{D_{emb}}\rightarrow\mathbb{R}^{2D_{emb}}$ is a two-layer perceptron with GELU activation, the $\textit{split}(\cdot,\text{dim}=2)$ operation divides the output features into two equal halves. This yields an approximation $\mathbf{A}_{\ell}\in\mathbb{R}^{H\times D_{emb}}$ and a detail $\mathbf{D}_{\ell}\in\mathbb{R}^{H\times D_{emb}}$. The $\mathbf{A}_{\ell}$ carries the smoother (low-frequency) information, while $\mathbf{D}_{\ell}$ captures the variations (high-frequency details) removed at level $\ell$. After $L$ levels, we obtain a final coarse approximation $\mathbf{A}_{L}$ and detail components $\{\mathbf{D}_{\ell}\}_{\ell=1}^{L}$. For convenience, we denote the set of multi-scale features as:
(10) $\mathcal{X}=\{\mathbf{X}^{(0)},\mathbf{X}^{(1)},\dots,\mathbf{X}^{(L)}\},$
where $\mathbf{X}^{(0)}:=\mathbf{A}_{L}$ represents the lowest-frequency band, and $\mathbf{X}^{(\ell)}:=\mathbf{D}_{L+1-\ell}$ for $\ell=1,\cdots,L$ (ordering from low to high frequency). Each $\mathbf{X}^{(\ell)}\in\mathbb{R}^{H\times D_{emb}}$ maintains the same dimensions as the original embedding.
@@
This multi-scale decomposition aligns with the established understanding in atmospheric science that different physical processes dominate at different scales (Hoskins et al., 1985), enabling our learnable approach to adaptively determine the optimal decomposition basis for S2S forecasting.
3.2.2. Physics-Informed Neural ODE